OpenStudy (anonymous):

Solve this system of equations using the addition method. x+y=6 -x+3y=-2 Solve this system of equations using the addition method. x+y=6 -x+3y=-2 @Mathematics

OpenStudy (hoblos):

y=1 x=5

OpenStudy (anonymous):

They're all lined up. Now just add through. x+y=6 -x+3y=-2 ----------- 0+4y=4 4y=4 divide both sides by 4. y=1 plug y back into the first equation (or any of the two for that matter) x+1=6 = x=5 or -x+3(1)=-2 = -x=-5 = x=5

OpenStudy (anonymous):

y=1 and x=5

OpenStudy (anonymous):

addinng both equations x+y=6 -x+3y=-2  4y=4 y=1 put y in any one x=5